OpenStudy (anonymous):

A square has an area of 16 cm². Calculate the length of its diagonal to the nearest 10th of a centimeter. . A. 6.4 cm B. 5.7 cm C. 8 cm D. 4 centimeters E. 64 cm

jhonyy9 (jhonyy9):

area = side * side =16 than one side = 4 cm diagonal = sqrt(16 +16) =sqrt32 =sqrt(16*2) =4sqrt2 = aprox. 5,7 B. is right

OpenStudy (anonymous):

Thanks that really helped

jhonyy9 (jhonyy9):

yw good luck bye